OpenStudy (anonymous):

m<10=4x; m<11=2x; x=___ HELP!

jimthompson5910 (jim_thompson5910):

are you given a pic of angles 10 and 11?

OpenStudy (anonymous):

I don't think so/:

jimthompson5910 (jim_thompson5910):

Then there's no way to know for sure how to connect angle 10 to angle 11

jimthompson5910 (jim_thompson5910):

Really odd how they just give you "m<10=4x; m<11=2x; x=___ " without a picture.

OpenStudy (anonymous):

There's a picture but I don't think it coincides with the problem

jimthompson5910 (jim_thompson5910):

go ahead and post it

OpenStudy (anonymous):

How do I do that?

jimthompson5910 (jim_thompson5910):

click "attach file"

jimthompson5910 (jim_thompson5910):

below the text box

OpenStudy (anonymous):

Oh its not on my computer..I'll just draw it for you

jimthompson5910 (jim_thompson5910):

ok that works

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

where are 10 and 11?

OpenStudy (anonymous):

That's what I was trying to say, I don't think there's a picture but It says they're complementary angles

jimthompson5910 (jim_thompson5910):

oh that helps then

jimthompson5910 (jim_thompson5910):

if angles 10 and 11 are complementary, then they add to 90 degrees

jimthompson5910 (jim_thompson5910):

So m<10 + m<11 = 90 4x + 2x = 90 6x = 90 x = 90/6 x = 15

OpenStudy (anonymous):

Ohhhhh I see!!! Thank you!

jimthompson5910 (jim_thompson5910):

np